OpenStudy (anonymous):

HELP PLEASE What is the solution of the system? y = 14x – 1 y = –11x + 4 A.(5, –51) B.(1/5,9/5) C.(5/3,67/3) D.(3/5,37/5) I would rlly apperciate help

OpenStudy (anonymous):

Make the two equations equal and solve for x

OpenStudy (anonymous):

How do I make them equal?

OpenStudy (anonymous):

y = y, so 14x - 1 = -11x + 4

OpenStudy (anonymous):

Im not getting it

OpenStudy (anonymous):

14x - 1 = -11x + 4, bring -11x over to give you 14x + 11x -1 = 4

OpenStudy (anonymous):

group the x and take the -1 over to give you 25x = 5

OpenStudy (anonymous):

solve for x?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

x = 5/25, simplify it to get 1/5

OpenStudy (anonymous):

sub into on of your equations to solve for the y

OpenStudy (anonymous):

K thx